Circular No. 8588

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 P3 (SWAN) Preliminary parabolic orbital elements from the CCD astrometry on

IAUC 8587 (predicted magnitudes from H = 11.0, n = 3):

T = 2005 Aug. 9.638 TT  ω = 30.^{o}616
Ω = 245.145 }2000.0
q = 0.51554 AU  i = 89.482
2005 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  Mag.
Aug. 23 11^{h}34.^{m}05 +36^{o}05.^{′}2 1.162 0.604 31.^{o}3 60.^{o}5  9.7
25 11 33.88 +37 46.5 1.181 0.629 32.2 58.8  9.9
27 11 33.41 +39 20.4 1.199 0.656 33.2 57.4  10.0
29 11 32.71 +40 47.6 1.215 0.685 34.3 56.2  10.2
31 11 31.82 +42 09.1 1.230 0.715 35.5 55.2  10.4
Sept. 2 11 30.79 +43 25.6 1.243 0.746 36.9 54.3  10.5
4 11 29.65 +44 38.0 1.254 0.777 38.3 53.5  10.7
6 11 28.43 +45 46.8 1.264 0.810 39.8 52.8  10.8
8 11 27.13 +46 52.8 1.272 0.842 41.4 52.2  11.0
10 11 25.78 +47 56.5 1.279 0.876 43.0 51.7  11.1
12 11 24.37 +48 58.4 1.285 0.909 44.7 51.2  11.2

Visual-magnitude and coma-diameter estimates: Aug. 26.14 UT, 10.5, 3′ (A. Hale, Cloudcroft, NM, 0.20-m reflector); 26.46, 9.0, 1′ (J. Kobayashi, Kumamoto, Japan, 20×100 binoculars).

SUPERNOVA 2005dl IN NGC 2276 A. Dimai, Cortina, Italy; and M. Migliardi, Tour Tour, France, report

discovery of an apparent supernova (mag 17.1) in the course of the CROSS program (cf. IAUC 7373) on four unfiltered CCD images taken on Aug. 25.1 UT with the 0.5-m ‘Ullrich’ telescope at Cortina d’Ampezzo, Italy. The new object, which was confirmed at mag 16.8 on three images taken on Aug. 26.13, is located at α = 7^{h}27^{m}31^{s}.75, δ = +85^{o}45′15′′.0 (equinox 2000.0), which is 18′′ east and 1′′ south of the nucleus of NGC 2276. Nothing is visible at this position on mages taken on Feb. 17 (limiting mag ∼ 19.0) by F. Manzini and on May 19 (limiting mag ∼ 18.5), and it is absent from Palomar Sky Survey red and blue plates.

2005 August 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green